OpenStudy (anonymous):

Four bells first begin to toll together and then at intervals of 6, 7, 8 and 9 seconds respectively. Find how many times the bells toll together in two hours and at what intervals they toll together?

OpenStudy (anonymous):

First find L.C.M of 6,7,8 and9

OpenStudy (anonymous):

ok. 504.

OpenStudy (anonymous):

Are u sure

OpenStudy (anonymous):

ya... thats right

OpenStudy (anonymous):

So, the four bells will ring afer 540 seconds all together

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

504.

OpenStudy (anonymous):

now, 2 hour = ---- seconds

OpenStudy (anonymous):

But it says they ring together at intervals of 6, 7, 8, 9 seconds.

OpenStudy (anonymous):

So what have these numbers got to do with 504?

OpenStudy (anonymous):

Since the Four bells first begin to toll together and then at intervals of 6, 7, 8 and 9 seconds respectively

OpenStudy (anonymous):

they will ring together after 540 seconds now

OpenStudy (anonymous):

PLZ read the question carefully

OpenStudy (anonymous):

So after that nine seconds, they will ring together after 504 seconds?

OpenStudy (anonymous):

nope.....

OpenStudy (anonymous):

Lets say they ring together first at 6.00.00. Then at 6.00.06 6.00.07 6.00.08 6.00.09 right?

OpenStudy (anonymous):

nope.....

OpenStudy (anonymous):

Let me first expalin the question

OpenStudy (anonymous):

there are four bells

OpenStudy (anonymous):

which ring together at first

OpenStudy (anonymous):

then BELL A rings after 6 seconds, BELL B rings after 7 seconds , BELL C rings after 8 seconds and BELL D rings after 9 seconds

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yep i got it now.

OpenStudy (anonymous):

so what do u think 540seconds means

OpenStudy (anonymous):

2 hours = 2 * 60 * 60 seconds = 3600 * 2 = 7200 seconds

OpenStudy (anonymous):

The three bells will ring TOGETHER every 504 seconds right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

OK, I know how to do it now. Thanks!!

OpenStudy (anonymous):

WELCOMX